OpenStudy (anonymous):

Explain two ways you could evaluate (f + g)(2).

OpenStudy (swissgirl):

Well either you can add the 2 functions and then plug in 2 and then solve Or You can plug in 2 and solve both equations separately and then add the 2 equations So it wld look like f(2)+g(2)

OpenStudy (swissgirl):

I will give you an example So let f(x)=1/x+3 and g(x)=x+2 #1 (f+g)(x)=(1/x+3)+(x+2)=1/x+x+5 Now lets plug in the 2 (f+g)(2)=1/2+2+5 = 15/2 #2 f(2)+g(2)=(1/(2)+3)+((2)+2) = (7/2)+(4) =15/2